40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Aug 19 11:30:20 2013, Jenne, Update, IOO, MC mirrors' ASC has non-zero inputs 
    Reply  Mon Aug 19 15:23:07 2013, Koji, Update, IOO, MC mirrors' ASC has non-zero inputs 
Message ID: 9032     Entry time: Mon Aug 19 15:23:07 2013     In reply to: 9030
Author: Koji 
Type: Update 
Category: IOO 
Subject: MC mirrors' ASC has non-zero inputs 

[Jenne, Koji]

This disturbance in the MC ASC channels were fixed.

This craziness happened ~10pm last night. Was there any action at the time? >> Sunday-night workers? (RXA: No, Nakano-kun and I left before 9:30 PM)

We found that the signals came from c1ioo. However, restarting, recompiling c1ioo and c1mcs didn't help
to clean up this issue. Just in case we cleaned up the corresponding entries in the ipc file /opt/rtcds/caltech/c1/chans/ipc/C1.ipc
and recomplied c1ioo and c1mcs because these are the channels we touched last week to mitigate the timing out issue of c1rfm.

Incidentally, we fell into a strange mode of the RCG: IOPs could not restart. We ended up running "sudo shutdown -r now"
on each machine (except for c1lsc which was not affected by this issue). This solved the issue.

Even now c1oaf could not be running properly. This is not affecting the IFO operation right now, but we need to look into this issue again
in order to utilize OAF.

ELOG V3.1.3-